OpenStudy (anonymous):

Simplify: n^2+3n+2 n+1 ----------- divided by ----- 2 2(n+2)

OpenStudy (ivettef365):

when you divide just switch the second equation and multiply so you have n^2 + 3n +2 2(n+2) ------------ . ----- ---- 2 n+1

OpenStudy (anonymous):

ok

OpenStudy (ivettef365):

solve for n^2+3n+2, do you know how to

OpenStudy (anonymous):

noo

OpenStudy (ivettef365):

that would be (n+2)(n+1) just because 2x 1 = 2 and 2+1 = 3 so you have (n+2)(n+1) 2(n+2) ---------- . ------- 2 (n+1)

OpenStudy (ivettef365):

now what can you convert to 1

OpenStudy (anonymous):

1 2 +2 -- . ------- 2 1 ?

OpenStudy (anonymous):

{(n^2+3n+2)/2}*{2(n+2)/(n+1)} ={(n^2+2n+n+2/2}*{2(n+2)/(n+1)} ={(n(n+2)+1(n+2)}*{2(n+2)/(n+1)} ={(n+1)(n+2)}*{2(n+2)/(n+1)} =2*{(n+2)^2}

OpenStudy (ivettef365):

you can cancel the (n+1) from top and bottom and you can cancel the "2" so you have n+2 n+2 ---- x ------ 1 1

OpenStudy (ivettef365):

now multiply and you have (n+2)^2 or you can write the answer as (n+2)(n+2) = n^2 +4n +4